OpenStudy (anonymous):

Suppose you work in a lab. You need a 15% acid solution for a certain test, but your supplier only ships a 10% solution and a 30% solution. Rather than pay the hefty surcharge to have the supplier make a 15% solution, you decide to mix 10% solution with 30% solution, to make your own 15% solution. You need 10 liters of the 15% acid solution. How many liters of 10% solution and 30% solution should you use?

OpenStudy (anonymous):

Can someone explain this to me step by step.

OpenStudy (tmak11):

im sorry the measurements i am trying to remember are to inaccurate and don't add up to the 15% A sol

OpenStudy (tmak11):

but it would seem you would need more of the 10% sol than the 30% sol

OpenStudy (anonymous):

Isn't there a method to solving this though

OpenStudy (tmak11):

yes..but i dont remember all the steps

OpenStudy (tmak11):

i think you take the 30 and use a 1 15 ratio to the 10

OpenStudy (tmak11):

to get the 15% needed

OpenStudy (jdoe0001):

\(\large { \begin{array}{rcccccl} &ltr(sol)&acidity&ltr(acid) \\\hline\\ 10\%&x&0.1&0.1x\\ 30\%\%&y&0.3&0.3y\\ mix&10&0.15&1.5 \end{array} \\ \quad \\\\ \quad \\ \quad \\ x+y=10\implies x={\color{brown}{ 10-y}}\qquad thus \\ \quad \\\\ \quad \\ \quad \\ \begin{array}{rcccccl} &ltr(sol)&acidity&ltr(acid) \\\hline\\ 10\%&{\color{brown}{ 10-y}}&0.1&{\color{blue}{ 0.1(10-y)}}\\ 30\%\%&y&0.3&{\color{blue}{ 0.3y}}\\ mix&10&0.15&{\color{blue}{ 1.5}} \end{array}\qquad thus \\ \quad \\ {\color{blue}{0.1(10-y)+0.3y=1.5 }} }\)